OpenStudy (anonymous):

Use the mthod of substitution to solv the linear equation. 2x-y=2 9x-4y=11

OpenStudy (anonymous):

2x-y=2......1 9x-4y=1.....2 from(1); 2x-2=y using the value of y in eq(2) 9x-4(2x-2)=1 9x-8x+4=1 x+3=0 x=-3........3 using value of xin eq(1) y=2(-3)-2 y=-6-2 y=-8

OpenStudy (anonymous):

ok i am going to see how u did this

OpenStudy (anonymous):

ok so could we do another probelm togther and i try to solve and u correct me?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

ok.....

OpenStudy (anonymous):

but i might need your help to start it off

OpenStudy (anonymous):

same method for linear equation 6a=5-b 6a=three-b

OpenStudy (anonymous):

my number three key is out...

OpenStudy (anonymous):

6a=5-b.......1 6a=3-b........2 from(1) 5-6a=b using it in two 6a=3-5+6a this question had no answer......

OpenStudy (anonymous):

i think i hav to set it up this way 5-6a=b

OpenStudy (anonymous):

my computer had frozee sorry for rplying late

OpenStudy (anonymous):

yeah...i have done it for you....

OpenStudy (anonymous):

no problem...

OpenStudy (anonymous):

ok so this problem is no solution

OpenStudy (anonymous):

wouldnt this problem be just a=6

OpenStudy (anonymous):

yeah......bcz variables cut down..

OpenStudy (anonymous):

because 5-three=1

OpenStudy (anonymous):

divide by 6=6

OpenStudy (anonymous):

oh ok i see i am going to graph this one and no solution. Thank you

OpenStudy (anonymous):

welcome....

OpenStudy (anonymous):

in which class you are???

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

intermediate

OpenStudy (anonymous):

what about u

OpenStudy (anonymous):

i am in university.......software engineering

OpenStudy (anonymous):

oh ok well im in healthscience

OpenStudy (anonymous):

well i am having a probelm with this particular homework assignment

OpenStudy (anonymous):

i dont remeber how to solve some of these problems